a_{2}^{0}(1320) ∗ ∗ ∗ ∗ ∗

March 25, 2009

The a_{2}^{0}(1320) is the neutral member of the isospin triplet from the lightest tensor multiplet. The mass and width are taken from [?]. The limit on the off-shellness of the particle is set to twice the width. The decay modes are similar to those in EvtGen apart from the modelling of ωππ as ωρ, the inclusion of the η′π and γγ modes the omission of the πγ mode. The a_{2}^{0}(1320) has mass 1318.3 MeV and is unstable. The a_{2}^{0}(1320) has spin 2, charge 0 and is colour neutral. The a_{2}^{0}(1320) is a meson and is from the 1^{3}P_{2} multiplet.The a_{2}^{0}(1320) has width 107 MeV. The lower limit on the mass of the particle is 214 MeV and the upper limit is 214 MeV. These are the deviations from the on-shell value. The branching ratios are fixed. The properties of the particle and its antiparticle are taken to be charge-conjugate to each other. The PDG code is 115.

Branching Rating On/ Outgoing  Description  Decayer
Ratio  Off Particles

0.349434 ∗ ∗ ∗ ∗ ∗ on ρ^{+}, π^{−} The decay of a_{2}^{0}(1320) to a rho and a pion, the branching ratio is from [?] with a minor change to ensure the branching ratios sum to unity. This is half the rate from [?] as there are two states.

Herwig::TensorMesonVectorPScalar::TensorVP

0.349434 ∗ ∗ ∗ ∗ ∗ on ρ^{−}, π^{+} The decay of a_{2}^{0}(1320) to a rho and a pion, the branching ratio is from [?] with a minor change to ensure the branching ratios sum to unity. This is half the rate from [?] as there are two states.

Herwig::TensorMesonVectorPScalar::TensorVP

0.144053 ∗ ∗ ∗ ∗ ∗ on π^{0}, η The decay of a_{2}^{0}(1320) to an η and a pion with branching ratio from [?] (there are minor adjustments to ensure the branching ratios sum to unity.)

Herwig::TensorMeson2PScalar::Tensor2PScalar

0.103473  ∗ ∗ ∗∗  on

ω, ρ^{0} The decay of a_{2}^{0}(1320) to two pions, modelled as the ρ, and an omega with branching ratio from [?] (there are minor adjustments to ensure the branching ratios sum to unity.)

Herwig::TensorMesonVectorVector::TVV

0.024184 ∗ ∗ ∗ ∗ ∗ on K^{+}, K^{−} The decay of the a_{2}^{0}(1320) to two kaons, this is half the rate from [?] due to the two possible final states. In addition there are minor modifications to ensure the branching ratios sum to unity.

Herwig::TensorMeson2PScalar::Tensor2PScalar

0.024184 ∗ ∗ ∗ ∗ ∗ on K^{0}, K̄^{0} The decay of the a_{2}^{0}(1320) to two kaons, this is half the rate from [?] due to the two possible final states. In addition there are minor modifications to ensure the branching ratios sum to unity.

Herwig::TensorMeson2PScalar::Tensor2PScalar

0.005229  ∗ ∗ ∗∗  on

η′, π^{0} The decay of a_{2}^{0}(1320) to a pion and an η′ with branching ratio from [?]. There are minor adjustments to ensure the branching ratios sum to unity.

Herwig::TensorMeson2PScalar::Tensor2PScalar

0.000009  ∗ ∗ ∗∗  on

γ, γ The decay of a_{2}^{0}(1320) to two photons, the branching ratio is from [?] with minor changes to ensure the branching ratios sum to one.

Herwig::TensorMesonVectorVector::TVV

Table 1: The decay modes of the a_{2}^{0}(1320).

The a_{2}^{0}(1320) decay modes are given in Table 1 and the total branching ratio is 1. The mass generator is a_20mass for the a_{2}^{0}(1320). The width generator is a_20width for the a_{2}^{0}(1320). The particle was checked by Peter Richardson at 15:59:57 on the 22nd of August 2007. The most recently changed decay mode

was modified at 15:20:45 on the 13th of October 2006. The particle data was last modified at 13:31:30 on the 21st of August 2007.